OpenStudy (anonymous):

What is the domain of the function: {(1, 2); (2, 4); (3, 6); (4, 8)}? A. {6, 8} B. {2, 4, 6, 8} C. {1, 2, 3, 4, 6, 8} D. {1, 2, 3, 4} What is the range of the function: {(2, 1); (4, 2); (6, 3); (8, 4)}? A. {2, 4, 6, 8} B. {1, 2, 3, 4} C. {1, 2, 3, 4, 6, 8} D. {6, 8} Suppose p varies directly with d, and p = 7 when d = 3. What is the value of d when p = 28? A.4/3 B.140/3 C. 16 D. 12

OpenStudy (danjs):

Domain - List all the inputs y=f(x) where x is the input Range - List all the resulting outputs, y=f(x) , y is the output in the set of points (x,y)

OpenStudy (danjs):

If p varies directly with d, then it is directly proportional , you can say... p = some constant x d p = k*d use given p and d, solve for the constant k value to get your expression

OpenStudy (anonymous):

ok i'll try that out

OpenStudy (anonymous):

where are you getting the k from?

OpenStudy (danjs):

it is a proportion constant, varies directly in general is y=k*x, linear

OpenStudy (danjs):

\[y~ \alpha~ x\] y=k*x